In a certain factory making razor blades there is a small probability of 1/500 for any blade
to be defective in each packet. The blades are supplied in packets of 10. Calculate the no.
of packets having a) no defective blade b) 1 defective blade c) 2 defective blade in a
consignment of 10,000 packets.
1
Expert's answer
2020-04-30T17:44:33-0400

p=0.002

L=10*0.002=0.02

P(n)=(Ln/n!)*e-L

a) P(0)=(0.020/0!)*e-0.02=0.98

N0=0.98*10000=9800

b)P(1)=(0.021/1!)*e-0.02=0.0196

N1=0.0196*10000=196

c)P(2)=(0.022/2!)*e-0.02=0.000196

N2=0.000196*10000 =2